What is the molarity (m) of a solution prepared by dissolving 29.25g NaCl in 3.00kg of water?

a. 0.0002
b. 0.17
c. 1.53
d. 9.75
1
Expert's answer
2020-04-18T06:59:37-0400

number of moles of "NaCl=\frac{29.25}{58.5}=0.5mol"

mass of solvent(water)"=3Kg"

molality, "m=\frac{moles of solute}{mass of solvent(in Kg)}=\frac{0.5}{3}=0.17molal"